Opened 9 years ago

Closed 3 years ago

#330 closed flight processing (fixed)

FW10/01, flight day 098/2010, New_Forest

Reported by: benj Owned by: knpa
Priority: alpha 4 medium Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2010/flight_data/uk/FW10_01-2010_098_New_Forest

Data arrived from ARSF via network transfer on 12th April 2010.

Scientific objective: Test flight for full-waveform LiDAR

Priority: a4m (not graded by SC, assigned following consultation with ppu)

PI: Ross Hill

Full waveform test flight. Will be delivered to PI but not specifically requested.

Sensors:

  • Eagle (18/08/2010)
  • Hawk (20/12/2010)
  • Leica LIDAR (Full Waveform) (11/08/2010)
  • RCD (11/08/2010)

Change History (39)

comment:1 Changed 9 years ago by mggr

Please notify Dr Hill when processing begins, as he and his student may wish to visit (once the data is confirmed to be basically ok).

comment:2 Changed 9 years ago by mggr

  • Summary changed from FW10_01, flight day 98/2010, Full waveform test to FW10/01, flight day 98/2010, Full waveform test

Fix summary line (in case we ever automate this).

comment:3 Changed 9 years ago by mark1

IPAS has been processed but due to large distance (~80km) to nearest available base station results are not great.

comment:4 Changed 9 years ago by mark1

applanix data also processed

comment:5 Changed 9 years ago by mark1

started hyperspectral processing

comment:6 Changed 9 years ago by mark1

  • Description modified (diff)
  • Summary changed from FW10/01, flight day 98/2010, Full waveform test to FW10/01, flight day 98/2010, New_Forest

comment:7 Changed 9 years ago by mark1

  • Owner set to mark1
  • Status changed from new to accepted

comment:8 Changed 9 years ago by mark1

Eagle file -6 has 3937 lines, not 1837 as reported in hdr file. Have edited hdr file appropriately.

comment:9 Changed 9 years ago by mark1

Removed photographs 1-16 since they do not overlap the flight lines

comment:10 Changed 9 years ago by mark1

Photograph delivery has been checked and looks ok.

comment:11 Changed 9 years ago by mark1

Hyperspectral delivery is ready for checking at:
/users/rsg/airborne/workspace/FW10_01-2010_098_New_Forest/delivery/20100524/FW10_01/

Lidar delivery is ready for checking (but needs an updated quality report) at:
/users/rsg/airborne/workspace/FW10_01-2010_098_New_Forest/delivery/20100525/FW10_01/

comment:12 Changed 9 years ago by mark1

Eagle and Hawk SCTs

Eagle Raw line SCT Hawk Raw Line SCT
-2 0.02 -2 -0.02
-3 0.04 -3 0.02
-4 -0.02 -4 -0.02
-5 0.03 -5 0.06
-6 -0.02 -6 0.01
-7 0.02 -7 0.01
-8 0.0 -8 0.03
-9 0.04 -9 0.01
-10 -0.04 -10 0.02
-11 0.02 -11 0.06
-12 0.0 -12 0.0
-13 0.0 -13 0.0
-14 0.0 -14 -0.01
-15 0.02 -15 0.02
-16 -0.02 -16 -0.01
-17 0.02 -17 0.06
-18 -0.02 -18 0.02
-19 0.02 -19 0.01
-20 0.0 -20 0.0
-21 0.0 -21 0.01
-22 0.0 -22 -0.02
-23 0.0 -23 0.01
-24 -0.02 -24 -0.02
-25 0.03 -25 0.02
-26 -0.04 -26 -0.02
-27 0.02 -27 0.0

comment:13 Changed 9 years ago by chrfi

delivery checked: all hawk lines, band 160 are value zero
no other issues

comment:14 Changed 9 years ago by mggr

For the full waveform delivery, please include ASCII + LAS1.0 versions of the discrete returns as well as LAS1.3 FW data. Also, we need to include the data sampling rate (possibly per line) in the documentation, e.g. 256 samples @ 1ns?

comment:15 Changed 9 years ago by emca

LIDAR data delivery checked.

Made some changes to the Readme and to the filenames to conform to the standard.
Checked several FW LAS files in TScan and all load okay with the .trj file.

Ready to deliver.

comment:16 Changed 9 years ago by mark1

  • Description modified (diff)

LiDAR and Photography data delivered to PI on 11/08/2010.

comment:17 Changed 9 years ago by mark1

PI requested that we send Eagle data separately now rather than wait for the Hawk calibration issue to be resolved.

comment:18 Changed 9 years ago by mark1

  • Description modified (diff)

Eagle data dispatched to PI 18/08/2010.

comment:19 Changed 9 years ago by mark1

There was no rinex/basestation data so used download option from grafnav when processing the trajectory data.

comment:20 Changed 9 years ago by mark1

Rsynced LiDAR/Camera/Eagle over to ~arsf. Did not rsync proc_laser since las files were delivered to PI in the LiDAR delivery.

comment:21 Changed 8 years ago by emca

Corrected LAS1.3 files sent to R Hill 21/10/2010

(incorrect header on original delivered files due to an error in processing software)

comment:22 Changed 8 years ago by anch

Delivery checked hawk - everything is fine and ready to deliver.

comment:23 Changed 8 years ago by mark1

  • Description modified (diff)

Hawk data has been dispatched to PI, as well as FW LiDAR extractions from the new extracting tool.

comment:24 Changed 8 years ago by emca

Hawk data re-delivered to PI 20/12/2010

comment:25 Changed 8 years ago by emca

  • Description modified (diff)

comment:26 Changed 8 years ago by mark1

  • Component changed from Processing: general to Archiving

rsynced remaining data over, except the proc_laser dir since las files were included with delivery.

ready to archive

comment:27 Changed 7 years ago by emca

Reprocessed hyperspectral data using APL software

No sync for hawk line 14, new sct value 1.45

Otherwise no problems

comment:28 Changed 7 years ago by knpa

  • Summary changed from FW10/01, flight day 98/2010, New_Forest to FW10/01, flight day 098/2010, New_Forest

comment:29 Changed 7 years ago by knpa

  • Owner changed from mark1 to knpa
  • Status changed from accepted to assigned

Beginning archiving.

comment:30 Changed 7 years ago by knpa

No basestation data.

comment:31 Changed 7 years ago by emca

Reprocessed data ready for delivery check

comment:32 Changed 7 years ago by knpa

Delivery checking the reprocessed hyperspectral.

comment:33 Changed 7 years ago by knpa

Delivery check complete

All looks good apart from:

  • Logsheet should have year in it's file name
  • Should mention regular dropped eagle scan lines in ReadMe

I've started zipping up the mapped files.

comment:34 Changed 7 years ago by emca

Renamed logsheet + edited readme

Ready for delivery once all mapped files have finished zipping

comment:35 Changed 7 years ago by emca

Data sent to R Hill 31/07/2012

Reprocessing complete - ready to be archived

comment:36 Changed 7 years ago by knpa

Made old deliveries available to NEODC 24/07/12

comment:37 Changed 6 years ago by knpa

Ingested by NEODC.

Moved to archive disk.

comment:38 Changed 6 years ago by knpa

Removed workspace version (original processing).

Leaving ticket open until reprocessing is archived.

comment:39 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

All data now available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/FW10_01/FW10_01-2010_098_New_Forest). Marking as closed.

Note: See TracTickets for help on using tickets.